# HARMONIC: Cognitive and Control Collaboration in Human-Robotic Teams

### v1

@@ Abstract
This paper presents a novel approach to multi-robot planning and collaboration. We demonstrate a cognitive strategy for robots in human-robot teams that incorporates metacognition, natural language communication, and explainability. The system is embodied using the HARMONIC architecture that flexibly integrates cognitive and control capabilities across the team. We evaluate our approach through simulation experiments involving a joint search task by a team of heterogeneous robots (a UGV and a drone) and a human. We detail the system’s handling of complex, real-world scenarios, effective action coordination between robots with different capabilities, and natural human-robot communication. This work demonstrates that the robots’ ability to reason about plans, goals, and attitudes, and to provide explanations for actions and decisions are essential prerequisites for realistic human-robot teaming.
@@ I Introduction
In today’s rapidly evolving technological landscape, robots are becoming increasingly ubiquitous across various domains, from manufacturing and healthcare to search and rescue operations [1, 2]. This proliferation has given rise to a growing number of heterogeneous robot teams, where machines with diverse capabilities collaborate to achieve common goals. To harness these distributed and varied problem-solving capabilities, research in Multi-Agent Planning (MAP) has focused on developing sophisticated planning and collaboration strategies.
Despite significant advancements in multi-robot planning and collaboration [3, 4], a number of issues related to human-robotic teaming remain relatively unexplored [5]. For robots to serve as trusted partners in human-robotic teams, they must demonstrate a number of cognitive and metacognitive abilities, including:
@@
Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning and low-level execution, respectively.
The above capabilities will facilitate complex team interactions, attention management, goal-setting, planning (including incorporating a variety of MAP strategies), and overcoming unexpected challenges in a way that humans can readily interpret. The integration of verbal and visual communication enhances the overall effectiveness of skill transfer and knowledge sharing within human-robot teams. Furthermore, meaning-oriented natural language communication allows robots to provide causal explanations for their actions and decisions, thereby enhancing overall transparency and explainability, which are stepping stones to trust [7]. A growing number of robots today are based on foundational models, such as Large Language Models (LLMs) [8, 9] and Vision-Language-Action (VLA) models [10, 11]. These black-box technologies contribute to several components of our architecture. However, we make sure that the modules whose operation and output should be explanatory to humans will be fully glass-box inspectable and explainable in human terms both to team members and to external inspectors.
We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task. We demonstrate the robots’ knowledge of team organization as well as, their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors. We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation.
Specifically, in this paper, we report:
@@
  • The distributed HARMONIC architecture, which enables a flexible integration of cognitive and robotic control and execution capabilities.
  • Results of simulation experiments on a search task carried out by heterogeneous robots, which showcases the human-robot team’s ability to handle complex, real-world scenarios.
Section II briefly describes the state of the art. Sections III-V present the HARMONIC framework and the implementation of its strategic and tactical levels. Section VI describes the organization of planning tasks, and Section VII presents our initial evaluation of HARMONIC on a team search task.
@@ II Background and Related Work
Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams encompasses a spectrum of strategies ranging from centralized to decentralized approaches, with varying degrees of inter-agent coupling. In centralized planning, a single agent orchestrates the entire team’s actions, potentially achieving globally optimal solutions but facing scalability challenges in larger teams. Conversely, decentralized methods enable autonomous plan negotiation among agents, offering enhanced scalability and adaptability at the risk of local suboptimality [12]. Brafman et al. [13] quantify this trade-off, noting that centralized planning in closely coupled multi-agent systems is exponentially more complex than solving independent problems in decoupled systems.
To balance these extremes, modern HMR teams often implement hybrid MAP strategies. These may include centralized hierarchical or collaborative planning coupled with decentralized execution [14], dynamic switching between local and global plans based on contextual factors [3], or centralized conflict resolution alongside decentralized planning [15]. Advanced implementations of these hybrid approaches include hierarchical planning frameworks, auction-based task allocation systems [16], and distributed partial-order planners [17], which have demonstrated significant collaborative capabilities in complex multi-robot scenarios.
Intra-team communication in an HMR team involves both inter-robot and human-robot interactions, each requiring distinct protocols. To date, most research has focused on inter-robot communication using standardized Agent Communication Languages to facilitate efficient information exchange [18]. The implementation of frameworks, such as CORA (Core Ontologies for Robotics and Automation) [19], while enhancing semantic interoperability and knowledge sharing among robotic agents, cannot support natural, human-level communication. To engender human trust, human-robot interactions must be carried out in everyday natural language.
The communication modalities currently employed in HMR teams, while effective for preprogrammed interactions between robots, Lemon et al. [20] and Natarajan et al. [5] support the need for a unified, human-centric approach to team communication. Rather than maintaining separate protocols for inter-robot and human-robot interactions, a common cognitive architecture that emulates human-like communication and reasoning offers a more cohesive solution, emulating human team dynamics.
Early work by Clark et al. [21] proposed a framework for this process, which was later adapted by Klein et al. [22] to team coordination, introducing the notion of common ground. These studies highlight essential capabilities for robotic team members, including maintaining shared understanding, negotiating goals and plans, managing attention, and handling commands and requests. Subsequent pioneering research in dialogue processing systems [23, 24, 25] and grounding [26, 27, 28, 29] further developed these ideas, laying the groundwork for addressing communication challenges in human-robot teams. Results of this research and our own prior work on team organization [30] and natural language communication [7] guided the implementation of cognitive processes for team organization and communication in HARMONIC.
III The HARMONIC Framework
HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig. 1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control. It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al. [31]. It also represents an advancement over the type 2 integration approach employed in the DIARC framework [32, 33]. Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
The strategic and tactical architectural layers of HARMONIC are connected by a bidirectional interface for seamless communication. The strategic layer, built upon the OntoAgent cognitive architecture [34, 35, 7], encompasses modules for attention management, perception interpretation, and decision-making (see Section. IV). It employs both utility-based and analogical reasoning, enhanced by metacognitive abilities. This layer prioritizes strategic goals, manages plan agendas, and selects actions while continuously monitoring their execution. It also facilitates team-oriented operations, including communicating in natural language and explaining decisions.
The tactical layer is responsible for robot control at a lower level, processing sensor inputs, and planning motor actions to execute high-level commands from the strategic layer. It employs controllers, algorithms, and models to translate abstract commands into concrete robot actions. For instance, a command PICK(key,at-position1) to ”pick up a key on the ground” initiates a series of precise operations, including object identification, position determination, trajectory computation, and actuation. This layer also manages reactive responses, such as collision avoidance, through dedicated controllers. In the tactical layer, HARMONIC employs Behavior Trees (BTs) [36], which provide modular and robust representations for skills and low-level plans (see Section. V).
The architecture’s design aligns with Kahneman’s dual-system approach [37], with the strategic layer implementing System 2 (slow, deliberative reasoning) and the tactical layer implementing System 1 (fast, intuitive responses). This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions. It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
IV OntoAgent for Strategic Level Planning
@@
OntoAgent is a content-centric cognitive architecture incorporated in HARMONIC. OntoAgent is designed to support the development of social intelligent agents through computational cognitive modeling [35, 7, 34]. This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action. The architecture’s memory structure is divided into three main components: a Situation Model (SM) that contains currently active concept instances; a Long-Term Semantic Memory (LTS) that stores knowledge about instances of events and objects, and an Episodic Memory (LTE) that stores knowledge about instances of events and objects. OntoAgent supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
@@
OntoAgent places a strong emphasis on natural language understanding and meaning-based text generation. These capabilities rely on a semantic lexicon that links words and phrases with meanings grounded in a resident ontology. The language analyzer treats a large range of complex linguistic phenomena (lexical disambiguation, reference resolution, ellipsis reconstruction, new-word learning, etc. [7]) and generates ontologically-grounded interpretations of text meaning. The language generator, for its part, generates natural language utterances from ontologically grounded meaning representations that the agent creates as part of its reasoning about action. OntoAgent also supports ontological interpretation of visual percepts using an opticon whose entries link images to ontological objects.
@@
The architecture supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
V Behavior Trees for Tactical Execution
@@
Behavior Trees (BTs), originally developed for controlling non-player characters in video games, have found significant applications in robotics and AI [36]. BTs are directed trees that provide a flexible and intuitive way to design control actions and define task-planning hierarchies for agents. The flexibility and modularity of BTs make them particularly suitable for complex, hierarchical task planning in robotics and AI applications, allowing for easy modification and scaling of agent behaviors.
@@
BTs in the tactical layer of HARMONIC enable effective reactive control and support the system’s safety and operational requirements in dynamic environments. The tactical layer also incorporates advanced control techniques such as whole-body compliant control [38, 36] with motion planning, path planning [39] etc. In addition to these, BTs can also be used for representing skills and low-level plans that can directly be shared between the robots[40, 41].
@@
BTs typically use a blackboard architecture to maintain and access frequently checked condition variables [42, 43]. In the context of HARMONIC, a state manager keeps track of these variables on the tactical layer and allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer.
Figure 2: (a) Root, control flow, and execution nodes in BTs. (b) BT design template for the tactical layer of robots running the HARMONIC architecture. (c) Sample BT on UGV.
BTs for HARMONIC robots follow a design template prioritizing safety and needs [40], as shown in Fig. 2. Collision avoidance subtrees are placed leftmost, followed by robot needs subtrees, then action command subtrees triggered by the strategic layer. Fallback subtrees on the far right provide actions when the robot is idle or other conditions aren’t met, such as random walks or waiting at the base station. This structure leverages the left-to-right execution of BTs to ensure proper prioritization of tasks.
The translation of the action commands to the appropriate state flags and control variables is done through the action data communication APIs, which modify the placeholder variables in the State Manager on the tactical layer.
VI Development of Distributed Plans
In this section, we show how robots implemented in HARMONIC carry out a search task. They decide to act based on reasoning carried out in the strategic layer. Their plans (which are instances of ontological scripts) are divided into steps that are communicated to the tactical layer by means of action commands.
VI-A High-level planning

### v2

@@ Abstract
This paper introduces HARMONIC, a cognitive-robotic architecture that integrates the OntoAgent cognitive framework with general-purpose robot control systems applied to human-robot teaming (HRT). We also present a cognitive strategy for robots that incorporates metacognition, natural language communication, and explainability capabilities required for collaborative partnerships in HRT. Through simulation experiments involving a joint search task performed by a heterogeneous team of a UGV, a drone, and a human operator, we demonstrate the system’s ability to coordinate actions between robots with heterogeneous capabilities, adapt to complex scenarios, and facilitate natural human-robot communication. Evaluation results show that robots using the OntoAgent architecture within the HARMONIC framework can reason about plans, goals, and team member attitudes while providing clear explanations for their decisions, which are essential prerequisites for realistic human-robot teaming.
@@ I Introduction
The proliferation of robots across various domains has given rise to a growing number of human-robot teams, where robots with various capabilities are employed by human operators to achieve complex mission goals [1]. Current state-of-the-art robotic systems demonstrate remarkable tactical proficiency, from the sub-millimeter precision of surgical robots performing intricate surgical procedures [2] to the dynamic multi-terrain locomotion of quadrupedal [3] and humanoid platforms exhibiting sophisticated whole-body coordination [4].
These systems also excel at integrating multi-modal sensory information through sensor fusion, simultaneously localizing themselves within complex environments, recognizing objects, tracking moving targets with increasing reliability, in adaptive policy learning and multi-robot planning and collaboration [5, 6].
Despite these significant advancements in tactical execution, most state-of-the-art robots are limited to short-horizon planning and execution and lack the strategic capabilities necessary for long-horizon planning, reasoning, and true collaborative partnership with humans in complex decision-making scenarios [7]. Importantly, effective human-robot teaming requires more than just tactical proficiency; to function as trusted partners within human-robot teams, they must exhibit a range of cognitive and metacognitive capabilities, including:
@@
Figure 1: An overview of the framework, showing the Strategic and Tactical components representing the high-level planning (System 2) and low-level execution (System 1), respectively.
Long-horizon functioning is within the scope of the field of cognitive architectures where much progress has been made in supporting complex agent team interactions, managing attention, setting goals, and conducting sophisticated planning, often incorporating diverse multi-agent planning strategies. Still, while cognitive architectures address key limitations in current robotic systems, they do not traditionally concentrate on the dynamic world of embodied agents. This has resulted in a notable disconnect between cognitive architectures and practical robotic control systems. Bridging this divide leverages the strengths of both these areas of research and development.
The HARMONIC framework presented in this paper directly addresses this challenge through the use of a dual-layer architecture that combines strategic cognitive capabilities with tactical robotic control. This integration not only enables long-horizon planning but also enhances the scope of the robot’s reasoning capabilities, adds metacognitive awareness and social intelligence, which in turn supports human-level explanation that serves as a foundation for robots to function as trusted teammates [10], rather than mere tools, within human-robot teams.
The cognitive architecture that we incorporate into HARMONIC as its strategic layer is OntoAgent [11, 12, 9]. It is responsible for the aforementioned cognitive architecture capabilities, thus facilitating complex team interactions, attention management, goal-setting, sophisticated planning (including the incorporation of diverse multi-agent planning strategies), and addressing unexpected challenges in ways that humans can readily interpret.
Additionally, OntoAgent integrates both verbal and visual communication, significantly enhancing skill transfer and knowledge sharing within human-robot teams, grounding, and supporting meaning-oriented natural language communication, enabling clear causal explanations for the robots’ actions and decisions. These capabilities greatly enhance transparency and explainability [10], establishing crucial foundations for trust in human-robot collaborations [9].
We present our approach to human-robot teaming using the example of a team comprised of two robots and a human executing a joint task. We demonstrate the robots’ knowledge of team organization as well as, their ability to (a) dynamically negotiate tasks and allocate responsibilities, (b) select and modify plans, and (c) adapt language communication to both the situation and the needs of their interlocutors. We implement these robotic capabilities using the novel HARMONIC architecture that operates in parallel at two levels – the strategic cognitive level, which supports operations requiring reasoning, and the tactical control level, which supports skill-based, ”automatic,” reflexive operation of robots.
Specifically, the key contributions of this paper are:
  • We introduce HARMONIC, a cognitive-robotic architecture, which enables a flexible integration of OntoAgent [11], a cognitive architecture with low-level robot planning and control using Behavior Trees [13].
@@
  • Results of simulation experiments on a search task carried out by a robotic team, which showcases the human-robot team’s interactions and ability to handle complex, real-world scenarios.
Section II briefly describes the state of the art and background. Section III introduces the HARMONIC framework and the implementation of its strategic and tactical levels. Section IV-B describes the task, the team configuration, and the development of distributed plans, and Section IV presents the evaluation.
@@ II Background and Related Work
II-A Cognitive-Robotic Architectures
There are relatively few cognitive architectures explicitly designed or effectively adapted for robotic applications due to the significant challenges involved in combining cognitive-level decision-making with motor control. Among these architectures, Soar [14], ACT-R [15], and DIARC [16] are prominent, each contributing distinct methods and philosophies toward cognitive autonomy in robotics.
The Soar architecture provides symbolic cognitive processing through rule-based decision-making, facilitating general problem-solving and reasoning [14]. It excels at clear, structured decision-making but requires extensive manual integration for perception and motor control, making it challenging to deploy in unpredictable or sensor-rich environments without substantial customization. The ACT-R architecture [15], on the other hand, models human cognition through symbolic reasoning combined with sub-symbolic activations, supporting sophisticated simulations of human cognitive processes like learning and memory. Its psychological plausibility makes it ideal for human-robot interaction scenarios, such as storytelling or social robotics, yet its heavy reliance on detailed knowledge engineering restricts flexibility in complex and dynamic environments.
Finally, DIARC is explicitly designed for cognitive robotics with distributed modules [16, 17], affective processes, and advanced natural language capabilities suitable for dynamic human-robot interactions. However, it struggles with concurrent process management, lacks clear architectural layering, unified ontology, explainability, long-term episodic memory, explicit metacognition, and formalized safety mechanisms, thus limiting its robustness and adaptability.
In summary, most architectures require extensive customization from both cognitive and control perspectives, leading to limited real-world applications and practical testing.
II-B Multi-Agent Planning (MAP) and Communication
Multi-Agent Planning (MAP) in Heterogeneous Multi-Robot (HMR) teams includes a variety of centralized and decentralized strategies, each presenting distinct trade-offs. Centralized planning achieves global optimality but encounters scalability challenges with larger teams, while decentralized methods improve scalability and adaptability but may result in suboptimal solutions [18], as highlighted by the exponential complexity identified by Brafman et al. [19]. Consequently, hybrid MAP approaches that blend centralized hierarchical planning with decentralized execution [20], dynamically switch between plans [5], or integrate centralized conflict resolution with decentralized planning [21] have gained traction. Advanced implementations of these hybrid strategies often incorporate auction-based task allocation systems and distributed planners [22, 23].
Regardless of the planning approach, effective intra-team communication remains integral for seamless coordination and execution within HMR teams. Such communication involves distinct protocols for human-robot and robot-robot interactions [24]. A unified, human-centric communication approach within a common cognitive architecture is advocated as essential by Lemon et al. [25] and Natarajan et al. [7]. Such architectures emulate human-like communication, reasoning, and team dynamics, building upon the notion of common ground introduced by Clark et al. [26] and further developed by Klein et al. [27]. Subsequent advances in dialogue processing systems [28] and grounding techniques [29, 30], together with insights from prior work on team organization [31] and natural language communication [9], guide current implementations of cognitive processes of OntoAgent in HARMONIC.
II-C The OntoAgent Cognitive Architecture
@@
OntoAgent is a content-centric cognitive architecture, designed to support the development of social intelligent agents through computational cognitive modeling [12, 9, 11]. This approach emphasizes the need to acquire, maintain, and dynamically expand large-scale knowledge bases, which are essential for an agent’s perception interpretation, reasoning, and action. The architecture’s memory structure is divided into three main components: a Situation Model (SM) that contains currently active concept instances; a Long-Term Semantic Memory (LTS) that stores knowledge about instances of events and objects, and an Episodic Memory (LTE) that stores knowledge about instances of events and objects. OntoAgent supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary.
@@
OntoAgent places a strong emphasis on natural language understanding (NLU) and meaning-based text generation. These capabilities rely on a semantic lexicon that links words and phrases with meanings grounded in a resident ontology. The language analyzer treats a large range of complex linguistic phenomena (lexical disambiguation, reference resolution, ellipsis reconstruction, new-word learning, etc. [9, 12]) and generates ontologically-grounded interpretations of text meaning. The language generator, for its part, generates natural language utterances from ontologically grounded meaning representations that the agent creates as part of its reasoning about action. OntoAgent also supports ontological interpretation of visual percepts using an opticon whose entries link images to ontological objects.
@@
The architecture supports goal-directed behavior through a goal agenda and a prioritizer that selects which goals to pursue. It typically uses stored plans associated with goals but can also engage in reasoning from first principles when necessary. In the current work, OntoAgent is incorporated in the strategic layer of HARMONIC.
II-D Behavior Trees
@@
Behavior Trees (BTs), originally developed for controlling non-player characters in video games, have found significant applications in robotics and AI [13]. BTs are directed trees that provide a flexible and intuitive way to design control actions and define task-planning hierarchies for agents. The structure of a BT includes control nodes (selectors, sequencers, and parallel nodes) that manage execution flow among child and execution nodes (action and condition nodes) that form the leaf nodes performing specific tasks. Control nodes process “ticks” at a set frequency, and conventionally execute from left to right, depending on the control node involved. This modular design enables easy modification and scaling of behaviors, making BTs particularly suitable for complex hierarchical task planning. In the current HARMONIC system, robot-specific BTs are integrated at the tactical layer to coordinate robotic control.
III The HARMONIC Framework
HARMONIC (Human-AI Robotic Team Member Operating with Natural Intelligence and Communication), shown in Fig. 1, is a dual-control cognitive-robotic architecture that integrates strategic, cognitive-level decision-making with tactical, skill-level robot control. It builds upon and extends the concept of hybrid control systems and architectures, as discussed in a comprehensive review by Dennis et al. [32]. It also represents an advancement over the type 2 integration approach employed in DIARC [16, 17]. Unlike DIARC, where the strategic layer is embedded as a subsystem within the tactical layer to enable concurrent and dynamic operations, HARMONIC introduces a more sophisticated integration of these components.
The strategic and tactical architectural layers of HARMONIC are connected by a bidirectional interface for seamless communication and data transfer. The strategic layer implements the OntoAgent cognitive architecture [11, 12, 9], encompasses modules for attention management, perception interpretation, and decision-making (see Sec. II-C). It employs both utility-based and analogical reasoning, enhanced by metacognitive abilities. This layer prioritizes strategic goals, manages plan agendas, and selects actions while continuously monitoring their execution. It also facilitates team-oriented operations, including communicating in natural language and explaining decisions.
The tactical layer is responsible for attention management, lower-level robot planning and control, sensor input processing, and executing motor actions based on high-level commands received from the strategic layer. The attention manager processes perception data, translating and prioritizing sensory information for use by the strategic component. Additionally, the tactical layer employs specialized controllers, algorithms, and models to convert abstract commands from OntoAgent in the strategic layer into precise, actionable robot operations.
For example, a command such as PICK(key, at-position1), from the OntoAgent to ”pick up a key on the ground,” triggers a well-defined sequence of tasks and actions including object identification, positional assessment, trajectory computation, and actuation, in the tactical layer. Furthermore, the tactical layer manages reactive behaviors, including collision avoidance, through dedicated controllers.
@@ III The HARMONIC Framework
The tactical layer of HARMONIC is built around the Behavior Trees (BTs) [13] which enable effective reactive control and support the system’s safety and operational requirements in dynamic environments. The tactical layer may also include advanced control policies for whole-body compliant control [33, 13] with motion planning, path planning [34] etc. In addition to these, BTs can also be used for representing skills and low-level plans that can also be shared between the robots [35, 36].
@@
In the context of HARMONIC, state manager (a blackboard) keeps track of condition and state variables on the tactical layer [37]. As shown in Fig. 2, this allows for efficient querying and updating of the system’s state during operation based on the sensory inputs and action commands received from the strategic layer.
Sensory inputs are continuously written into the State Manager (SM), where the attention module filters and packages this perception data for the OntoAgent in the strategic layer. The high-level action commands sent by the OntoAgent through Interface APIs are unpacked for updating the state variables in the SM. While sensor data and physical state tracking operate in real time, the strategic layer runs asynchronously. Real-time control is achieved through the data available in the SM that is accessed by the LiveBT in parallel, which chooses policies and actions stored in the Action Manager. This architecture enables HARMONIC to maintain real-time behavior while supporting higher-level planning capabilities.
Figure 2: Data flow and control diagram showing the connections between the BTs in the tactical layer and the OntoAgent in the strategic layer in the HARMONIC System.
Action commands from the strategic layer are translated into corresponding state and control variables via action data communication APIs. These APIs update the placeholder variables within the State Manager, ensuring accurate and timely control of the tactical layer operations.
The architecture’s design aligns with Kahneman’s dual-system approach [38], with the OntoAgent in the strategic layer implementing System 2 (slow, deliberative reasoning) and the BTs in the tactical layer implementing System 1 (fast, intuitive responses). This structure allows for dynamic scheduling and adaptation, enabling real-time adjustments to priorities and actions. It effectively handles computational delays, contingencies, safety concerns, and resource optimization through a combination of low-level planning and reactive algorithms.
IV Task Design
We evaluated the HARMONIC framework’s teaming strategy using a simulated search task set in an apartment environment, as illustrated in Fig. 4.6. The scenario involves an HMR team consisting of an unmanned ground vehicle (UGV), a drone, and a human team member. The objective is to find a set of lost keys. The human initiates the task by remotely communicating with the robotic team members.
The simulation environment, developed using the Unity game engine, incorporates abstracted versions of robot behaviors. This approach allowed us to focus on key aspects of the HARMONIC framework’s performance while simplifying the implementation of complex robot actions that are not central to our current work.
IV-A Heterogenous Robot Team Configuration
The simulation evaluation involved two robots: an Unmanned Ground Vehicle (UGV) and a drone, both functionally and structurally heterogeneous, offering complementary capabilities and perspectives [39]. Each robot is equipped with sensors for localization, collision detection, and object recognition, facilitating coordinated and efficient operations.
The UGV’s design (see Fig. 3.a) enables it to access ground-level areas and maneuver under obstacles and surfaces. The drone, on the other hand, effectively complements the UGV by performing aerial search operations. The drone’s (see Fig. 3.b) vertical mobility and sensing capabilities allow it to scan environments extensively from above, navigating over obstacles and accessing areas difficult for the UGV. Since the robots communicate in natural language, their dialogue and actions are comprehensible to humans, and humans can intervene when necessary.
Both robots operate decentralized instances of tactical and strategic modules, ensuring efficient and autonomous team planning and control. Currently, the UGV acts as the team leader at the task level, responsible for managing interactions with the human operator and coordinating with the drone. The leader assignment is presently randomized; however, future work aims to develop and evaluate heuristic methods for dynamic team hierarchies. This advancement will optimize leader selection based on contextual factors and the individual capabilities of each agent relevant to specific tasks.
IV-B Development of Distributed Plans
Plans are distributed between strategic and tactical components in both UGV and drone systems, with each robot acting on reasoning from the OntoAgent in the strategic layer. The high-level plans (ontological script instances) are divided into steps communicated to the tactical layer as action commands, which are executed through low-level control encoded in BTs.
IV-B1 High-level planning